Draw and label a rectangle that has a perimeter of 60cm. and an area of 200 square centimeters?

Answer by ewatrrr(24785)

 
Hi,

p = 2*l + 2*w
60 = 2*l + 2*w
30 = l + w OR w = 30-l
A = w*l = (30-l)*l = 200cm^2
l^2 -30l +200 = 0
(l-20)(l-10)= 0
(l-20)= 0 l = 20 OR
(l-10)= 0 l = 10
rectangle 10 by 20 satisfies the conditons